Baseball. A catcher catches a 145 g fastball traveling horizontally at 42.5 m/s. How large an impulse does the ball give to catcher? If the ball takes 20.0 ms to stop once it is in contact with the catcher’s mitt, what average force did the ball exert on the catcher? What is the baseball's average acceleration during this interval? (answer is in meters per square second, you only need to type the number not the units on this one)
Answer #1

A)

Impulse, J = mv

J = 0.145 x 42.5 = 6.16 kg m/s

B)

Impulse = Ft = mv

Force, F = 6.16/0.02 = 308.13 N

Acceleration, a = F/m = 308.13/0.145 = 2125 m/s^2

Comment in case any doubt please rate my answer....